OpenStudy (anonymous):

find the sum (8x2 + 17x) + 6x2

Parth (parthkohli):

\[\begin{aligned}8x^2 + 17x + 6x^2 &=& 8x^2 + 6x^2 + 17x \\ & = & (8 + 6)x^2 + 17x \end{aligned}\]

OpenStudy (anonymous):

thank you

Parth (parthkohli):

What might \(8 + 6\) be?

OpenStudy (anonymous):

14

Parth (parthkohli):

Yes, so that becomes \(14x^2 + 17x\).

OpenStudy (anonymous):

ok i understand now thanks